OpenStudy (anonymous):

can someone give me answers for these two questions please :) Question 2 Solve for r in the equation C = 2d – 2r A) r = C + 2d – 2 B) r = -c + 2d/2 C) r = c + 2d/ -2 D) r = C + 2d Question 4 A) (0, –3) B) (0, 5) C) (0, 3) D) (0, –5)

OpenStudy (anonymous):

the question for #4 is What is the y-intercept of the graph of 3x – 5y = 15

OpenStudy (anonymous):

C = 2d – 2r -2d from both sides. -2d+C=-2r -2r=C+2d --- ------ -2 -2 r=C/-2 -d

OpenStudy (anonymous):

Thats for question 2 ^^^

OpenStudy (anonymous):

For question 4 you just plot out the graph, and the y intercept is -3.

OpenStudy (anonymous):

so its 0,-3

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

ok wait so #2 which answer choice is it is it a,b,c,or d?